OpenStudy (anonymous):

-(m-8)=3m-4(1-2m) how do you do this with a explaination?

OpenStudy (anonymous):

-(m-8) = 3m - 4(1-2m) Start by break out the parentheses -m + 8 = 3m -4*1 + -4 * -2m -m + 8 = 3m -4 + 8m -m + 8 = 11m - 4 Adding the m's 8 = 11m + m - 4 = 12m + 4 Moving over the -m 8 + 4 = 12m 12 = 12m 12/12 = 12m / 12 1 = m divides both sides by 12 m = 1

OpenStudy (anonymous):

thanks!